OpenStudy (anonymous):

I know this may sound like a dumb question but I am having a memory lapse...how do you find the derivative of (1)/(2X)

OpenStudy (anonymous):

\[= (2x)^{-1}\] use the chain rule and power rule

OpenStudy (anonymous):

I am just kidding the problem was (1)/(x^2-1)

OpenStudy (anonymous):

or\[=\frac{ 1 }{ 2 }x ^{-1}\] just the power rule

OpenStudy (anonymous):

ok thank you very much:)

OpenStudy (anonymous):

then rewrite like i did in the first post and use the chain rule

OpenStudy (anonymous):

\[\frac{ 1 }{ x^{2}-1 }=(x^{2}-1)^{-1}\]

OpenStudy (anonymous):

right so then it would be -1(x^2-1)^-2 times (2x)

OpenStudy (anonymous):

\[\frac{ d }{ dx }(x^{2}-1)^{-1} = -(2x)(x^{2}-1)^{-2}=-\frac{ 2x }{ (x^{2}-1)^{2} }\]

OpenStudy (anonymous):

\[-1(x^{2}-1)^{-2}(2x)\]

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

ok thanks \[\frac{ 2x }{ x ^{4}-2x+1 }\]

OpenStudy (anonymous):

oh with a negative in front?

OpenStudy (anonymous):

why did you multiply that out? it really isn't necessary and could be a pain if you were to take further derivatives.

OpenStudy (anonymous):

because i am actually using it to find the points of the graph here the tangent line is horizontal so i have to set it equal to 0 and solve for the point

OpenStudy (anonymous):

you'll end up factoring it again. the bottom is 0 when x = 1 or -1. however, the tangent will not be horizontal there but rather it will be vertical. set the top = 0 and solve to find horizontal tangents.

OpenStudy (anonymous):

ohhh lol you're right

OpenStudy (anonymous):

sooo im confused as to what to do after setting it equal to 0

OpenStudy (anonymous):

solve for x. that's where you'll get a horizontal tangent.

OpenStudy (anonymous):

i know that but how

OpenStudy (anonymous):

\[-\frac{ 2x }{ (x^{2}-1)^{2} }=0 \Rightarrow 2x = 0 \text{ solve}\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ohh so set both the top and the bottom equal to 0?

OpenStudy (anonymous):

\[2x=0 and (x ^{2}-1)^{2}=0\]

OpenStudy (anonymous):